OpenStudy (anonymous):

What are the solutions of 2x2 + 8x = -26?

OpenStudy (anonymous):

-2 ± 3i -2 ± i 2 ± 6i -3 ± 2i

OpenStudy (anonymous):

POP, POP!

OpenStudy (anonymous):

2(x^2 + 4x) = -26 x^2 + 4x = -13

OpenStudy (anonymous):

Now we use (b/2)^2 which is 4 and add that to both sides: x^2 + 4x + 4 = -9 This is (x+2)^2 = -9 rad both sides x+2 = rad(-9) x + 2 = i * rad(9) x + 2 = +-3i x = -2 +- 3i